OpenStudy (anonymous):

Find the least common multiple: 3x^2, 12y, and 10x^3y^3.

OpenStudy (anonymous):

6x^2y

OpenStudy (anonymous):

Are we sure about the answer 6x^2y?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Ohh wait wait is 12y 2nd term or a part of 1st term???

OpenStudy (anonymous):

Hello

myininaya (myininaya):

3x^2=3*x*x 12y=2*2*3*y 10x^3y^3=2*5*x*x*x*y*y*y the last one has the most y's and x's the middle one has no x's we need x^3 there the top one has no y's we need y^3 there the last one has 5 in it both, we need 5 in the other products the middle one has most twos, we had the least amout of 2's in the first one we need 2*2 there and we a 3 in the last once since we have a 3 in the other prime factorization so that is 3*2*2*5*y^3*x^3=60y^3x^3

myininaya (myininaya):

I hope my thinking made sense.

OpenStudy (anonymous):

3x^2 = 1 * 3 * x * x 12y = 2 * 2*3*y 10x^3y^3 = 2* 2 *5 * x * x *x *y *y*y 2*2*3*5*x*x*x*y*y*y = 60x^3y^3

myininaya (myininaya):

i like to go in z y w ... c b a order lol @ Doodle

OpenStudy (anonymous):

or check this out: The lcm of 3, 12, and 10 is 60 Get the highest expoent of x, which is x^3 Get the highest expeonet of y which is y^3 This our answer is 60x^3y^3

OpenStudy (anonymous):

If 12 y is another term then the L.C.M is 3

OpenStudy (anonymous):

To find the LCM follow these steps (1) Write the prime factorization of each number (2) List each prime factor the Greatest number of times it appears in any factorization (3) Multiply the prime factors in the list

OpenStudy (anonymous):

LCM of 12, 10 and 3 is 3??

myininaya (myininaya):

lcm(12,10,3) is not 3

myininaya (myininaya):

lcm(12,10,3)=2*2*3*5

OpenStudy (anonymous):

Of course not, you guys are making this more complicated than it should be

myininaya (myininaya):

I like my process but it sounds funny to me.

OpenStudy (anonymous):

Thats casue it is, and its super long

myininaya (myininaya):

lol

OpenStudy (anonymous):

This is a person who is learning so it is best for her/him to learn a process and then once the concept is understood he or she can take short cuts and do it in their head or whatever. The process for doing these is below and I demonstrated in my example. To find the LCM follow these steps (1) Write the prime factorization of each number (2) List each prime factor the Greatest number of times it appears in any factorization (3) Multiply the prime factors in the list